OpenStudy (surryyy):

Write the equation y=-1/2x+3 in standard form.

OpenStudy (freckles):

well standard form looks like this Ax+By=C And some teachers prefer A,B,C to be integers So this means we need to get that term with an x on the side with the term with the y

OpenStudy (freckles):

That is one step to perform.

OpenStudy (freckles):

There is just one more after that.

imqwerty (imqwerty):

standard form of a straight line= ax+by+c=0 y=(-1/2)x+3 multiply by -2 throughout and then bring it to to that form :)

OpenStudy (surryyy):

1/2x + y = -1/2x + 1/2x + 3 y + 1/2x = 3 like that?

OpenStudy (freckles):

yes and you can write that as 1/2*x+y=3 now to get all the coefficients to be integer multiply by 2 on both sides (this will get rid of that fraction thingy)

OpenStudy (freckles):

@surryyy are you having problems?

imqwerty (imqwerty):

multiplying 2 on both sides - 2(1/2x +y) =2(3) x+2y=6 x+2y-6=0

OpenStudy (freckles):

I would probably leave as x+2y=6

OpenStudy (freckles):

I think Ax+By+C=0 this is general form but sometimes called standard form

OpenStudy (surryyy):

oh ok ty

imqwerty (imqwerty):

:)